### v2

# On the Learnability of Watermarks for Language Models

Abstract
Watermarking of language model outputs enables statistical detection of model-generated text, which has many applications in the responsible deployment of language models. Existing watermarking strategies operate by altering the decoder of an existing language model, and the ability for a language model to directly learn to generate the watermark would have significant implications for the real-world deployment of watermarks. First, learned watermarks could be used to build open models that naturally generate watermarked text, allowing for open models to benefit from watermarking. Second, if watermarking is used to determine the provenance of generated text, an adversary can hurt the reputation of a victim model by spoofing its watermark and generating damaging watermarked text. To investigate the learnability of watermarks, we propose watermark distillation, which trains a student model to behave like a teacher model that uses decoding-based watermarking. We test our approach on three distinct decoding-based watermarking strategies and various hyperparameter settings, finding that models can learn to generate watermarked text with high detectability. We also find limitations to learnability, including the loss of watermarking capabilities under fine-tuning on normal text and high sample complexity when learning low-distortion watermarks. [^1] We publicly release code and trained model weights at https://github.com/chenchenygu/watermark-learnability.
1 Introduction
As large language models (LLMs) become more capable and widely used, watermarking language model outputs becomes increasingly important to mitigate potential harms and misuses of LLMs. Watermarking enables statistical detection of model-generated text, which enables enforcing policies around LLM usage, e.g., removing LLM-generated disinformation from social media platforms or detecting academic dishonesty. Another proposed use case of watermarking is identifying text provenance, i.e., tracing a text back to the specific LLM that generated it (Abdelnabi & Fritz, 2021; Kuditipudi et al., 2023).
Recent works have demonstrated that it is possible for an LLM provider to inject specific, known watermark patterns into text using specialized decoding algorithms (Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023), but little is known about whether these watermarks are learnable by a model. The learnability of watermarks has significant implications for the real-world deployment of watermarks, as it could enable downstream applications and adversarial attacks.
In this work, we study the learnability of watermarks by studying weights-based watermarking, which involves learning parameters for a language model that cause it to generate watermarked text under its natural sampling distribution, without using a special decoding-time watermark algorithm. Our investigation is motivated by its relevant implications for two applications: developing watermarking for open language models and the possibility of watermark spoofing attacks.
Existing watermarking methods depend upon using a specialized decoding algorithm. This makes them too brittle for open models, as users will use a variety of decoding algorithms due to user choice, different inference infrastructures and environments, etc. Weights-based watermarking removes the reliance on a specialized decoder and is robust to changes in decoding algorithm, making it a necessary step towards building open models that generate watermarked text using only their weights.
Watermark spoofing attacks are hypothetical attacks proposed in Sadasivan et al. (2023) where an adversary outputs text that contains the watermark signal from a victim model. If watermarking is used to identify text provenance, then such an attack can attribute harmful behavior to and damage the reputation of the victim model. The ability to learn watermarks from samples of watermarked text raises the possibility of spoofing attacks. Successful execution of spoofing attacks suggests limitations for using watermarks to identify provenance, which has implications for the responsible usage and deployment of watermarking.
Figure 1: A diagram illustrating decoding-based watermarking (top) vs. weights-based watermarking (bottom). Notice that decoding-based watermarking requires a specialized decoding algorithm $f_{\text{w}}$ to generate watermarked text, whereas weights-based watermarking generates watermarked text directly from the model, using just its weights. Watermark distillation enables weights-based watermarking by training a student model $p_{\theta}$ to behave like the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$.
To enable weights-based watermarking, we propose logits-based and sampling-based watermark distillation, two simple methods for a student model to learn weights-based watermarking from a teacher model with decoding-based watermarking. Intuitively, in logits-based watermark distillation, the student model is trained to match the next token distribution outputted by the teacher model using decoding-based watermarking. In sampling-based watermark distillation, the teacher model with decoding-based watermarking is first used to generate watermarked samples. Then, the student model is fine-tuned on these watermarked samples.
We experiment with three distinct decoding-based watermarking strategies: KGW (Kirchenbauer et al., 2023a), Aar (Aaronson, 2023), and KTH (Kuditipudi et al., 2023), and various values for their hyperparameters. We find that watermarks and hyperparameter settings vary in their degree of learnability. For example, in each watermarking strategy, higher-distortion hyperparameter settings are successfully learned by both forms of watermark distillation (median p-values less than 0.0001). Lower-distortion watermarks and hyperparameter settings are more challenging and less sample efficient to learn, but not unlearnable, as the p-values are still noticeably smaller than random chance.
We demonstrate applications of these findings towards developing watermarks for open-source models and the possibility of spoofing attacks. We find that weights-based watermarking is robust to changes in decoding algorithm and is a promising first step towards developing watermarking for open models. However, these watermarking capabilities are removed by further fine-tuning on normal text, indicating that improving robustness to fine-tuning is an important remaining challenge. For spoofing, we show that sampling-based watermark distillation is applicable to the spoofing setting, raising the possibility of spoofing attacks. We demonstrate a proof-of-concept attack on instruction-following chat models to substantiate this possibility, which suggests limitations on using watermarks to assign provenance and blame to specific models.
2 Background and notation: decoding-based watermarking
We study autoregressive language models $p_{\mathrm{LM}}:\mathcal{V}^{\ast}\to\Delta(\mathcal{V})$ that map from a prefix string $x$ to a next token distribution over the vocabulary $\mathcal{V}$.
Informally, a decoding-based watermarking strategy uses a watermark key to modify the model’s original next token distribution into a new distribution for generating watermarked text, which has a watermark signal embedded. The watermark detection algorithm looks for this signal using the same watermark key. Formally, we define a decoding-based watermarking strategy to be a function
$f_{\text{w}}:\Delta(\mathcal{V})\times\mathcal{V}^{\ast}\times\Xi\to\Delta(\mathcal{V})$ (1)
where $\Xi$ is the set of possible watermark keys. This function $f_{\text{w}}$ outputs a distribution $p_{\text{w}}(\cdot\mid x)$ from which to generate the next token in the watermarked text, given an original next token distribution $p_{\mathrm{LM}}(\cdot\mid x)$, input text $x$, and watermark key $\xi\in\Xi$.
We define a watermark detection algorithm to be a function
$f_{\text{d}}:\mathcal{V}^{\ast}\times\Xi\to[0,1]\,.$ (2)
Given some text $x$ and watermark key $\xi$, $f_{\text{d}}$ outputs a p-value with respect to the null hypothesis that $x$ is independent of $f_{\text{w}}$ with key $\xi$. Internally, $f_{\text{d}}$ computes a test statistic that measures the strength of the signal embedded by the watermark, then computes a p-value using the distribution of the test statistic under the null hypothesis. If the p-value is below a certain significance level, the null hypothesis is rejected and the text is detected as watermarked. Slightly imprecisely, rejecting the null means the text is detected as model-generated. [^2] This is slightly imprecise because model-generated text is not the only text that can be watermarked. For example, someone could write watermarked text by manually following the watermarking algorithm. However, for most practical use cases, such as detecting academic dishonesty, this minor imprecision is not an issue because either way, the user is doing something suspicious and unusual.
In this paper, we use three decoding-based watermarking strategies: Algorithm 2 in Kirchenbauer et al. (2023a), the Gumbel softmax scheme in Aaronson (2023), and the exponential minimum sampling scheme in Kuditipudi et al. (2023). Using the authors’ names and initials, we refer to these as KGW, Aar, and KTH, respectively.
2.1 KGW
In the KGW watermarking strategy (Kirchenbauer et al., 2023a), when generating the next token, the vocabulary is pseudorandomly split into a “green list” and “red list” by hashing the previous token. The green list contains watermark hyperparameter $\gamma\in(0,1)$ proportion of the vocabulary. Hyperparameter $\delta>0$ is added to the original log-probabilities of the green list tokens, which are then softmaxed. This procedure makes green list tokens more likely in watermarked text than in non-watermarked text. At detection time, if the proportion of green list tokens in a text is much greater than $\gamma$, then the p-value is small.
Formally, the KGW decoding-based watermarking strategy is defined as
$f_{\text{w}}^{\text{KGW}}\left({\bm{p}},x,\xi;\gamma,\delta\right)=\mathrm{softmax}\bigl{(}\log({\bm{p}})+\delta\cdot f_{\text{hash}}^{\text{KGW}}\left(x_{\mathrm{len}(x)};\xi,\gamma,|\mathcal{V}|\right)\bigr{)}$ (3)
where $f_{\text{hash}}^{\text{KGW}}$ is parameterized by key $\xi$ and hashes the previous token $x_{\mathrm{len}(x)}$ and returns $g\in\{0,1\}^{|\mathcal{V}|}$, which contains $\gamma\cdot|\mathcal{V}|$ ones and $(1-\gamma)\cdot|\mathcal{V}|$ zeros, encoding the green list. The KGW watermark detection function is
$f_{\text{d}}^{\text{KGW}}\left(x,\xi;\gamma\right)=1-F_{B}\left(\sum\nolimits_{t=2}^{\mathrm{len}(x)}f_{\text{hash}}^{\text{KGW}}\left(x_{t-1};\xi,\gamma,|\mathcal{V}|\right)_{x_{t}}\right)$ (4)
where $F_{B}$ is the cumulative distribution function (CDF) for binomial distributed random variable $B\sim\mathrm{Bin}(\mathrm{len}(x)-1,\gamma)$. $f_{\text{w}}^{\text{KGW}}$ distorts model outputs by upweighting green list tokens, increasing perplexity of generated texts computed by a large model (Kirchenbauer et al., 2023a). Increasing the bias hyperparameter $\delta$ increases detectability, i.e., smaller p-values, but also increases distortion.
2.2 Aar
The Aar watermarking strategy (Aaronson, 2023) uses a pseudorandom hash function $f_{\text{hash}}^{\text{Aar}}$ parametrized by key $\xi$ that hashes the previous $k$ tokens and returns $r\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$. $r$ assigns a score to each token in the vocabulary, and $k\in\mathbb{N}$ is a hyperparameter. The Aar decoding-based watermarking strategy is defined as
$f_{\text{w}}^{\text{Aar}}\left({\bm{p}},x,\xi;k\right)=\mathrm{onehot}\left(\operatorname*{arg\,max}_{i}\,f_{\text{hash}}^{\text{Aar}}\Bigl{(}x_{\mathrm{len}(x)-k+1},\ldots,x_{\mathrm{len}(x)};\xi,|\mathcal{V}|\Bigr{)}_{i}^{1/{\bm{p}}_{i}},|\mathcal{V}|\right)$ (5)
where $\mathrm{onehot}(i,|\mathcal{V}|)$ returns a $|\mathcal{V}|$-dimensional probability vector with $1$ at index $i$ and 00 everywhere else. Aar chooses the token that maximizes $r_{i}^{1/p_{i}}$, i.e., a token $i$ with both a high original probability $p_{i}$ and hash score $r_{i}$. The Aar watermark detection function is
$f_{\text{d}}^{\text{Aar}}\left(x,\xi;k\right)=1-F_{G}\left(\sum\nolimits_{t=k+1}^{\mathrm{len}(x)}-\log\Bigl{(}1-f_{\text{hash}}^{\text{Aar}}\left(x_{t-k},\ldots,x_{t-1};\xi,|\mathcal{V}|\right)_{x_{t}}\Bigr{)}\right)$ (6)
where $F_{G}$ is the CDF for gamma distributed random variable $G\sim\mathrm{Gamma}(\mathrm{len}(x)-k,1)$. Because $f_{\text{w}}^{\text{Aar}}$ is deterministic based on the previous $k$ tokens and the original model probabilities, Aar can sometimes produce repetitive text, especially at small $k$ (Kuditipudi et al., 2023). Increasing $k$ decreases repetition as $k$-grams become less likely to be repeated, but the watermark also becomes less robust to edits, as each token edit affects the hashed watermark scores for $k+1$ tokens.
2.3 KTH
The KTH watermarking strategy (Kuditipudi et al., 2023) is similar to Aar, but instead of hashing the previous tokens to obtain the watermark scores $r$, the scores are obtained from the next position in the key sequence $\xi$. In KTH, $\xi=\xi^{(1)},\ldots,\xi^{(m)}$ where each $\xi^{(j)}\in[0,1]^{|\mathcal{V}|}$ with entries uniformly distributed across $[0,1]$ and $m$ should be longer than the maximum generation length. Then, the KTH decoding-based watermarking strategy is defined as
$f_{\text{w}}^{\text{KTH}}\left({\bm{p}},x,\xi\right)=\mathrm{onehot}\left(\operatorname*{arg\,max}_{i}\,\left(\xi^{(\mathrm{len}(x))}_{i}\right)^{1/{\bm{p}}_{i}},|\mathcal{V}|\right)$ (7)
where $\mathrm{onehot}$ is as defined for Aar. Note that $f_{\text{w}}^{\text{KTH}}$ is deterministic. To allow the model to generate different responses to the same prompt, before generating each sequence, $\xi$ can be shifted by some random $\tau$, i.e., $\xi^{\prime}=\xi^{(1+\tau\bmod m)},\ldots,\xi^{(m+\tau\bmod m)}$ and $\xi^{\prime}$ is used for $f_{\text{w}}^{\text{KTH}}$ (Kuditipudi et al., 2023). To study the impact of these shifts on learnability, we introduce an additional hyperparameter $s\in[1,m]$ which restricts the range of the shift (e.g., $s=1$ means the original $\xi$ is always used, and $s=m$ means that any shift of $\xi$ can be used). [^3] We space the $s$ shifts evenly across $[1,m]$, e.g., the set of possible shifts $\tau$ is $\{i\cdot\lfloor m/s\rfloor:0\leq i<s\}$. Increasing $s$ expands the range of possible generations the model can output.
At detection time, the basic test statistic is $\phi(x,\xi)=\sum_{t=1}^{\mathrm{len}(x)}-\log(1-\xi^{(t)}_{x_{t}})$. However, to be robust to edits, the detection statistic computes a minimum Levenshtein distance using the basic test statistic as a cost (see Definition 5 in Kuditipudi et al. (2023)) and compares it to the expected distribution under the null (see Algorithms 5 and 6 in Kuditipudi et al. (2023)). [^4] This expected reference distribution has finite number of samples $T$, and the p-values obtained by comparing to this distribution are lower bounded by $\frac{1}{T+1}$.
3 Methods
3.1 Problem statement
Given a teacher language model $p_{\mathrm{LM}}$, decoding-based watermarking strategy $f_{\text{w}}$, corresponding detection algorithm $f_{\text{d}}$, and key $\xi$, the goal is to learn a student model $p_{\theta}$ whose sampling distribution naturally generates watermarked text. If $p_{\theta}$ generates text that attains small detection p-value $f_{\text{d}}(y,\xi)$ with probability similar to that of $p_{\mathrm{LM}}$ with $f_{\text{w}}$, then we say that $p_{\theta}$ has learned a weights-based watermarking strategy, since $p_{\theta}$ has learned to generate watermarked text using just its weights. See Figure 1 for an illustration of decoding-based versus weights-based watermarking.
Next, we present two methods for learning a weights-based watermarking strategy: logits-based watermark distillation and sampling-based distillation, which fall under the broader category of knowledge distillation (Hinton et al., 2015; Kim & Rush, 2016).
3.2 Logits-based watermark distillation
Informally, in logits-based watermark distillation (WD-L), we train the student model to behave as if it had a decoding-based watermarking strategy applied. We train the student model $p_{\theta}$ to match the next token distribution outputted by the teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ and key $\xi$, i.e., $f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x_{<t}\right),x_{<t},\xi\right)$. Formally, the training objective is to minimize the KL divergence loss
$\mathcal{L}_{\text{WD-L}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{x\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(x)}D_{\mathrm{KL}}\left(f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid x_{<t}\right),x_{<t},\xi\right)\,\|\,p_{\theta}\left(\cdot\mid x_{<t}\right)\right)$ (8)
where $\mathcal{D}$ is the training dataset and teacher model $p_{\mathrm{LM}}$ is frozen. This approach requires that $p_{\mathrm{LM}}$ and $p_{\theta}$ have the same tokenizer and vocabulary so that the logits can be aligned between the two models. It is also helpful if both $p_{\mathrm{LM}}$ and $p_{\theta}$ have the same model architecture, as in this case we can initialize $p_{\theta}$ to $p_{\mathrm{LM}}$. One advantage of this approach is that the ground truth next tokens $x_{t}$ from dataset $\mathcal{D}$ is not used in the loss function, and thus $\mathcal{D}$ does not need to be watermarked text; standard datasets containing non-watermarked human-generated text can be used.
3.3 Sampling-based watermark distillation
Sampling-based watermark distillation (WD-S) consists of two stages. First, we generate watermarked text from teacher model $p_{\mathrm{LM}}$ with decoding-based watermarking strategy $f_{\text{w}}$ applied using key $\xi$. Then, we fine-tune the student model $p_{\theta}$ on this watermarked text using the standard language modeling cross-entropy loss.
Formally, given a prompt $x$, we generate a watermarked completion sequence $y=y_{n}y_{n-1}\cdots y_{1}$ where $y_{t}\sim f_{\text{w}}\left(p_{\mathrm{LM}}\left(\cdot\mid y_{<t}x\right),y_{<t}x,\xi\right)$. Given a set of prompts $\mathcal{P}$, we populate a finetuning dataset $\mathcal{D}$ with watermarked sequences $y$ from prompts $x\in\mathcal{P}$. Then, we train $p_{\theta}$ to minimize the cross-entropy loss
$\mathcal{L}_{\text{WD-S}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{y\in\mathcal{D}}\sum_{t=1}^{\mathrm{len}(y)}-\log p_{\theta}\left(y_{t}\mid y_{<t}\right).$ (9)
Here, $p_{\mathrm{LM}}$ and $p_{\theta}$ do not need to share the same tokenizer or vocabulary. However, sampling-based watermark distillation is less efficient than logits-based watermark distillation due to autoregressively generating watermarked text.
4 Experimental setup
4.1 Watermarking strategies
We experiment with all three decoding-based watermarking strategies discussed in §2. We use various hyperparameter settings to vary the level of distortion induced by the watermarks. Specifically, we test KGW with $\gamma=0.25$ and $\delta=\{1,2\}$, [^5] Because we always use $\gamma=0.25$, we sometimes omit explicitly stating the value of $\gamma$ to reduce clutter. Aar with $k=\{2,3,4\}$, and KTH with key length 256 and number of shifts $s=\{1,2,4,256\}$.
4.2 Training
For each watermarking strategy, we test logits-based and sampling-based watermark distillation.
For logits-based watermark distillation, we use Llama 2 7B (Touvron et al., 2023) as both the teacher and student models (the student model is initialized with the teacher model weights). We distill using a subset of OpenWebText (Gokaslan et al., 2019) for 5,000 steps with a batch size of 64, sequence length of 512 tokens, [^6] For KTH we use a batch size of 128 and sequence length of 256 tokens because we use key length 256. maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix A.1.
For sampling-based watermark distillation, we use Llama 2 7B as the teacher model and Pythia 1.4B (Biderman et al., 2023) as the student model. [^7] We include additional experiments using Llama 2 7B as both the teacher and student model for sampling-based watermark distillation in Appendix C. Note that Llama 2 and Pythia have different tokenizers, so sampling-based distillation is necessary for this setting. First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640k watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. [^8] We include additional experiments using fewer training samples to investigate sample efficiency in Appendix D. Then, we fine-tune Pythia 1.4B on the watermarked samples for 1 epoch, roughly 8,000 steps, with a batch size of 64, sequence length of 256 tokens, [^9] Note that Llama 2 and Pythia tokenize sequences differently. Pythia has a larger vocabulary size and tends to tokenize sequences into fewer tokens compared to Llama 2. maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup. Full training details are in Appendix A.2.
4.3 Evaluation and metrics
Evaluation procedure. As in Kirchenbauer et al. (2023a) and Kuditipudi et al. (2023), we evaluate on generations prompted by prefixes from the RealNewsLike subset of the C4 dataset (Raffel et al., 2020). For each decoding-based watermarking strategy and distilled model, we generate 5k 200-token completions from 50-token prompts from the validation split. We use standard sampling with temperature 1 for the main results, and investigate the model’s robustness to different decoding parameters in §5.3. We include evaluations on additional datasets in Appendix F.
We choose metrics to evaluate two properties: watermark detectability and generation quality.
Watermark detectability. We compute the median watermark detection p-value across generations. [^10] For watermark detection, we truncate to the first 200 tokens using the detection tokenizer (Llama 2). Note that the p-values for the KTH watermark are lower bounded by how many samples $T$ we compute in the reference distribution. Similar to Kuditipudi et al. (2023), we use $T=10,000$, [^11] As in Kuditipudi et al. (2023), we do not use higher $T$ due to the slowness of computing the test statistic. so the p-values are lower bounded by 1e-04. So, to make fine-grained distinctions in watermark strength below this lower bound, we also compute the median test statistic (defined in §2.3) to evaluate KTH watermark strength. A larger test statistic indicates higher watermark detectability.
We also compute the AUROC (area under the receiver operating characteristic curve) for classifying model-generated versus human-generated text using the watermark detection p-values/test statistics. We compute the AUROC using an equal number of model-generated and human-generated texts, all of the same length.
Generation quality. We compute the mean LM-score across generations, which we define to be perplexity evaluated by Llama 2 13B. [^12] Note that LM-score does not perfectly capture quality or fluency, as repetitive text has a low LM-score, so we also use separate metrics measuring repetitiveness. We also compute MAUVE to measure distributional similarity between model generations and human text (Pillutla et al., 2021). To evaluate repetitiveness, we compute mean seq-rep-3 across generations, which is the proportion of duplicate 3-grams in a sequence (Welleck et al., 2020).
Comparisons and notation. For both watermark distillation experiments, for each decoding-based watermarking strategy $f_{\text{w}}$, we compare the teacher model with $f_{\text{w}}$ (denoted by DecT) against the distilled student model (denoted by WD-L and WD-S for logits-based and sampling-based watermark distillation, respectively). As a baseline for generation quality, we use the base student model with no watermarking or distillation (denoted by “Base student”).
For the sampling-based watermark distillation experiment, since the teacher Llama 2 7B model is larger and more powerful than the student Pythia 1.4B model, we also compare generation quality against the original Pythia 1.4B model with $f_{\text{w}}$ (denoted by DecS) to control for model size. This allows for a fair comparison of generation quality between decoding-based watermarking and sampling-based watermark distillation.
5 Results
[…]
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 4.4e-18 6.9e-09 1.000 0.997 16.4 17.5 0.04 0.03 0.89 0.90
$\delta=1$ 1.0e-05 6.0e-03 0.992 0.913 12.6 13.1 0.03 0.03 0.91 0.90
Aar $k=2$ 1.3e-75 2.2e-12 0.999 0.996 6.5 10.8 0.34 0.12 0.28 0.83
$k=3$ 4.8e-73 1.1e-01 1.000 0.773 9.5 11.5 0.14 0.04 0.74 0.93
$k=4$ 4.0e-72 3.7e-01 0.999 0.580 10.7 11.8 0.09 0.03 0.86 0.92
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (569) 0.994 0.998 10.5 16.6 0.03 0.04 0.88 0.83
$s=2$ 1.0e-04 (600) 1.0e-04 (478) 0.998 0.994 10.7 16.5 0.03 0.04 0.91 0.84
$s=4$ 1.0e-04 (593) 1.0e-03 (439) 0.995 0.969 10.5 14.0 0.03 0.04 0.91 0.90
$s=256$ 1.0e-04 (600) 4.6e-02 (426) 0.996 0.893 10.6 11.4 0.03 0.04 0.92 0.92
Base student 5.0e-01 0.500 11.7 0.03 0.92
Table 1: Results for the logits-based watermark distillation experiments. Within each watermark type, the hyperparameter rows go from higher-distortion to lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently.
Figure 2: Empirical cumulative distribution functions (eCDFs) of the distributions of watermark detection p-values on generations from logits-based (left) and sampling-based (right) watermark distilled models on various watermark types and hyperparameter settings. In higher-distortion watermarks, the majority of generations have small p-values. In lower-distortion watermarks, the p-values are larger, but still consistently smaller than a non-watermarked baseline.
5.1 Logits-based watermark distillation
Table 1 contains results for logits-based watermark distillation. [^13] As discussed in §4.3, we report the median KTH test statistic in parentheses to allow for finer-grained distinctions in watermark strength below the p-value lower bound of 1.0e-04. We find that logits-based watermark distillation (WD-L) successfully learns higher-distortion [^14] Here, we are using “distortion” somewhat informally, roughly meaning how much of a difference watermarking causes in terms of generation quality, behavior, etc. watermarks, achieving small p-values and high detectability. Lower-distortion watermarks are harder to learn, but they are still learnable, just less efficiently. Table 1 shows that WD-L does not achieve as small watermark detection p-values as decoding-based watermarking, but for higher-distortion watermark hyperparameter settings (larger $\delta$ for KGW, smaller $k$ for Aar, and smaller $s$ for KTH), the p-values are still sufficiently small to enable high detectability, as shown by the high AUROC scores. Figure 2 (left) contains empirical CDFs of the distributions of p-values across generations, showing that for higher-distortion watermarks, the majority of generations from WD-L have small p-values.
Within each watermark type, WD-L p-values are larger for lower-distortion hyperparameter settings, indicating that lower-distortion watermarks are harder to learn. These watermarks are still learned to some degree, however, as the p-values are noticeably smaller than the random chance baseline of 0.5, and the AUROC scores are noticeably higher than the baseline of 0.5. We also find that if we run logits-based watermark distillation on Aar $k=3$ for five times longer (25,000 steps) on more data, the median p-value decreases from 0.11 to 0.00014. This suggests that lower-distortion watermarks are less sample efficient to learn, but still learnable, given sufficient training data and steps.
In addition, we find that for each watermark type, WD-L weights-based and decoding-based watermarking distort generation quality along the same axes, although the degree of distortion varies. For example, for Aar, both WD-L and decoding-based watermarking are more repetitive than base Llama 2 7B, especially at lower $k$. However, WD-L induces less distortion than decoding-based Aar, as shown by the lower repetition and higher MAUVE scores of WD-L. On the other hand, WD-L induces more distortion than decoding-based watermarking on KGW and KTH, as shown by the higher LM-scores of WD-L.
5.2 Sampling-based watermark distillation
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 4.4e-18 3.6e-06 1.000 0.995 16.4 32.8 51.9 0.89 0.81 0.65
$\delta=1$ 1.0e-04 2.0e-02 0.992 0.874 12.6 26.2 41.3 0.91 0.84 0.72
Aar $k=2$ 1.3e-75 2.9e-14 0.999 0.993 6.5 7.0 19.8 0.28 0.10 0.52
$k=3$ 4.8e-73 1.6e-02 1.000 0.866 9.5 15.1 29.0 0.74 0.46 0.68
$k=4$ 4.0e-72 3.0e-01 0.999 0.631 10.7 20.5 33.8 0.86 0.76 0.75
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (456) 0.994 0.989 10.5 22.2 34.0 0.88 0.72 0.81
$s=2$ 1.0e-04 (600) 4.0e-04 (445) 0.998 0.972 10.7 21.9 26.3 0.91 0.79 0.84
$s=4$ 1.0e-04 (593) 2.2e-03 (436) 0.995 0.955 10.5 22.2 24.7 0.91 0.84 0.85
$s=256$ 1.0e-04 (600) 2.1e-03 (437) 0.996 0.955 10.6 22.2 25.7 0.92 0.85 0.88
Base student 5.0e-01 0.500 21.7 0.88
Table 2: Results for the sampling-based watermark distillation experiments. Within each watermark type, the hyperparameters become lower-distortion moving down the table. Higher-distortion watermarks are successfully learned with small p-values and high detectability. Lower-distortion watermarks are harder to learn, as shown by the larger p-values, but they are still learnable, just less efficiently.
Table 2 contains results for the sampling-based watermark distillation experiments. [^15] As discussed in §4.3, we report the median KTH test statistic in parentheses to allow for finer-grained distinctions in watermark strength below the p-value lower bound of 1.0e-04.161616The seq-rep-3 metric is deferred to Appendix B due to space constraints. The trend is the same as in logits-based watermark distillation, with Aar increasing repetitiveness across all methods, especially at low $k$. We find the same trends and conclusions as in the logits-based watermark distillation experiments. Table 2 and Figure 2 (right) show that sampling-based watermark distillation (WD-S) successfully learns higher-distortion watermarks, achieving small p-values and higher detectability. The p-values from WD-S are not as small as those from decoding-based watermarking, but still small enough to enable high detectability, as shown by the high AUROC scores. Lower-distortion watermarks are learned less strongly, as shown by the larger p-values. However, in sample complexity experiments in Appendix D, we find that more training samples and steps lead to smaller p-values, with no sign of convergence at 640k samples, which is how many we use here. This suggests that lower-distortion watermarks are less sample efficient to learn, but still learnable, given enough training samples.
5.3 Robustness to changes in decoding algorithm
nucleus sampling temperature sampling
Model $p=1$ $p=0.95$ $p=0.9$ $p=0.85$ $t=0.75$ $t=0.5$ $t=0.25$ $t=0$
KGW $\delta=2$ 6.9e-09 2.9e-09 1.2e-09 4.8e-10 1.9e-10 2.8e-11 3.7e-12 1.3e-12
Aar $k=2$ 2.2e-12 4.6e-13 8.5e-14 3.9e-14 5.2e-15 9.9e-17 9.9e-18 5.3e-18
KTH $s=1$ 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04 1.0e-04
Table 3: Watermark detection p-values of generations from logits-based watermark distilled Llama 2 7B models under various decoding algorithms and parameters. We observe small p-values in all settings, showing that weights-based watermarking is robust to changes in the decoding algorithm.
Whereas decoding-based watermarking relies on special decoding algorithms, weights-based watermarking generates watermarked text naturally, i.e., under standard decoding algorithms. In Table 3, we compute watermark detection p-values from weights-based watermarking under a variety of decoding algorithms and parameters: nucleus sampling (Holtzman et al., 2020) with different $p$, temperature-based sampling with different temperatures $t$, and greedy decoding ($t=0$).
We find that weights-based watermarking is robust to changes in the decoding algorithm, achieving consistently high detectability. We also observe that p-values decrease as $t$ or $p$ decreases, with greedy decoding obtaining the smallest p-values.
5.4 Robustness to text edits
We test the robustness of weights-based watermarking to edits by randomly corrupting generated text from logits-based watermark distilled Llama 2 7B models with varying proportions of tokens randomly edited. See Appendix G for full experimental details. As shown in Figure 4, all three watermarks are robust to moderate proportions of tokens edited, up to around 0.2-0.3. The watermark remains detectable with small p-values. At higher proportions of random edits, up to around 0.6, KTH is significantly more robust to edits than KGW and Aar, consistent with the findings of Kuditipudi et al. (2023).
Figure 3: Watermark detection p-values on generations from weights-based watermarking, corrupted with varying proportions of tokens randomly edited. The watermarks are robust to moderate amounts of corruption.
Figure 3: Watermark detection p-values on generations from weights-based watermarking, corrupted with varying proportions of tokens randomly edited. The watermarks are robust to moderate amounts of corruption.
Figure 4: Watermark detection p-values of generations from logits-based watermark distilled Llama 2 7B models after further fine-tuning on OpenWebText. The models’ weights-based watermarking is removed by fine-tuning.
6 Applications
6.1 Watermarking for open models
In §5.3, we showed that weights-based watermarking is robust to changes in the decoding algorithm. This is a necessary first step towards watermarking for open models, where users can run inference themselves. They may change the decoding algorithm, and the inference library they use may not enable decoding-based watermarking by default or implement it at all.
Robust watermarking for open models should also ideally be robust to fine-tuning, as users have the ability and desire to fine-tune open models for their own use cases. However, Figure 4 shows that weights-based watermarking obtained from watermark distillation is not robust to further fine-tuning on normal, non-watermarked text (see Appendix H for experimental details). We leave addressing this challenge and learning weights-based watermarking that is robust to fine-tuning to future work.
However, weights-based watermarking also has potential use cases that do not require robustness to further fine-tuning. For example, weights-based watermarking could be used for watermarking open models which are unlikely to be fine-tuned further by users, such as RLHF-trained instruction-following chat models. In addition, weights-based watermarking simplifies decoding compared to decoding-based watermarking, as there is no need for an additional specialized decoding algorithm. So, weights-based watermarking can easily be deployed into existing highly optimized infrastructures and inference algorithms, as it just requires loading different model weights.
6.2 Spoofing attacks
Abdelnabi & Fritz (2021) and Kuditipudi et al. (2023) consider the possibility of using watermark detection to attribute the provenance of generated text to a specific model, which could help policy enforcement and auditing of model providers. However, using watermarking for provenance attribution brings the risk of spoofing attacks: an adversary generates damaging text containing the watermark of a victim model, making it appear as if the victim model generated it, thus hurting the reputation of the victim model (Sadasivan et al., 2023). Sampling-based watermark distillation is applicable to the spoofing attack setting, as only black-box samples from the victim/teacher model are needed for distillation.
In this proof-of-concept experiment, we simulate a spoofing attack using a victim model of Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$). Llama 2-Chat 7B is trained for safety and tends to refuse harmful requests (Touvron et al., 2023). The goal of the spoofing attack is generate watermarked responses to harmful requests, damaging the victim model’s reputation for safety. We obtain an adversary model by performing sampling-based watermark distillation with Alpaca-7B (Taori et al., 2023) as the student and the victim model Llama 2-Chat 7B as the teacher. We query the victim model for watermarked samples, filter out refusals, then fine-tune the adversary model on those samples. See Appendix I.1 for full training details.
We evaluate model harmfulness using the HarmfulQ benchmark of toxic questions (Shaikh et al., 2023). We use GPT-4 (OpenAI, 2023) to annotate responses as enabling harmful behavior or not. See Appendix I.2 for full evaluation details. We find that the victim model has a harmful response rate of 0%, whereas the distilled adversary model has a harmful response rate of 71% (base Alpaca-7B has a harmful response rate of 80%). Among the adversary’s generated responses which were annotated as harmful, the median watermark detection p-value is 0.002 (with a median generation length of 593 tokens), [^17] Among all 200-token slices from each of the harmful responses, the median detection p-value is 0.04. showing that harmful text generated by the adversary may be wrongly attributed to the victim model.
7 Related work
Post-hoc detection. Many works have studied post-hoc detection of model-generated text, without modifying the generation process itself. Some works train a binary classifier to perform detection (Zellers et al., 2019; Bakhtin et al., 2019; Tan et al., 2020), see Jawahar et al. (2020) for a survey. Other methods are zero-shot, using heuristics and metrics for detection (Gehrmann et al., 2019; Solaiman et al., 2019; Mitchell et al., 2023). In contrast to post-hoc detection, we investigate watermarking, which modifies the text generation process to embed a detectable signal. However, post-hoc detection could potentially be used in conjunction with watermarking (Mitchell et al., 2023).
Text watermarking. Older works on textual watermarking edit pre-existing text to inject signals that can be statistically detected (Rizzo et al., 2019; Abdelnabi & Fritz, 2021; Yang et al., 2022), see Kamaruddin et al. (2018) for a survey. More specifically, we study decoding-based watermarking which modify decoding procedures to generate new watermarked text (Venugopal et al., 2011; Kirchenbauer et al., 2023a; Aaronson, 2023; Kuditipudi et al., 2023; Zhao et al., 2023a; Christ et al., 2023; Hu et al., 2023; Wu et al., 2023). Recent works have proposed various watermarking algorithms, including semantic watermarks (Fu et al., 2023; Hou et al., 2023; Liu et al., 2023b; Ren et al., 2023), multi-bit watermarking (Yoo et al., 2023; Wang et al., 2023), and public/private key watermarking (Liu et al., 2023a; Fairoze et al., 2023), but note that these are also all decoding-based. Decoding-based watermarking is not robust to changes in decoding algorithm, whereas we find that weights-based watermarking is.
Watermark attacks. Prior works have investigated attacks to remove the watermark from watermarked text, e.g., by paraphrasing, swapping tokens, etc. (Kirchenbauer et al., 2023b; Krishna et al., 2023; Sadasivan et al., 2023; Zhang et al., 2023). In addition, watermark spoofing attacks are where an adversary generates watermarked text that falsely appears as if it came from a victim model. Sadasivan et al. (2023) investigates some spoofing attacks, but they require knowledge of how the watermark signal works. Our spoofing attack using sampling-based watermark distillation does not requre such knowledge.
Watermarking for protection against model extraction. A number of prior works have studied watermarking for protection against model extraction/stealing attacks, where an adversary “steals” a model via API access by distilling on its outputs (He et al., 2022a; Zhao et al., 2022; He et al., 2022b; Zhao et al., 2023b). In this setting, the suspect model is queried multiple times on specially chosen inputs to detect whether the model was stolen. The watermark is specifically designed to be learnable by the adversary model via distillation. On the other hand, text watermarking enables detecting whether any text was model-generated from any input prompt, so the object of detection is different (model versus text). Existing decoding-based text watermarking strategies are designed to be subtle and imperceptible in order to minimally impact text quality. They are not specifically designed to be learnable, so their learnability must be determined via empirical investigation.
8 Conclusion
In this paper, we investigate the learnability of watermarks for language models. We find that logits-based and sampling-based watermark distillation are able to successfully learn weights-based watermarking that is robust to changes in decoding algorithm and text edits, although lower-distortion watermarks are harder and less sample efficient to learn. Our findings address a key technical challenge towards developing watermarking for open models and raise the possibility of watermark spoofing attacks against victim models.
Future work may explore improving robustness of weights-based watermarking to further fine-tuning, which would address another important challenge towards robust watermarking for open models. Future work may also more comprehensively study and evaluate spoofing attacks and potential defenses against spoofing attacks, which would have implications for whether watermarks should be used to assign provenance and blame to specific models.
Acknowledgments
We gratefully acknowledge the support of an Open Philanthropy Project Award. Chenchen Gu was supported by a Stanford CURIS Fellowship. Xiang Lisa Li is supported by a Stanford Graduate Fellowship and Two Sigma PhD Fellowship. Tatsunori Hashimoto is supported by a gift from Open Philanthropy and by the Tianqiao and Chrissy Chen Institute.
Appendix A Watermark distillation training details
A.1 Logits-based watermark distillation training details
We train using a subset of OpenWebText for 5,000 steps with a batch size of 64, sequence length of 512 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 6 hours on 4 NVIDIA A100 80GB GPUs.
A.2 Sampling-based watermark distillation training details
We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch, roughly 8,000 steps, with a batch size of 64, sequence length of 256 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Each training run took approximately 3 hours on 1 NVIDIA A100 80GB GPU.
Appendix B Full sampling-based watermark distillation results table
Table 4 contains the full version of Table 2, including values for the seq-rep-3 metric, which was deferred here due to space constraints. We observe the same trends for seq-rep-3 as in logits-based watermark distillation (Table 1), where Aar increases repetition in all methods, especially for lower $k$, whereas KGW and KTH do not affect repetition.
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 4.4e-18 3.6e-06 1.000 0.995 16.4 32.8 51.9 0.04 0.04 0.02 0.89 0.81 0.65
$\delta=1$ 1.0e-04 2.0e-02 0.992 0.874 12.6 26.2 41.3 0.03 0.03 0.02 0.91 0.84 0.72
Aar $k=2$ 1.3e-75 2.9e-14 0.999 0.993 6.5 7.0 19.8 0.34 0.53 0.23 0.28 0.10 0.52
$k=3$ 4.8e-73 1.6e-02 1.000 0.866 9.5 15.1 29.0 0.14 0.27 0.10 0.74 0.46 0.68
$k=4$ 4.0e-72 3.0e-01 0.999 0.631 10.7 20.5 33.8 0.09 0.12 0.05 0.86 0.76 0.75
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (456) 0.994 0.989 10.5 22.2 34.0 0.03 0.03 0.02 0.88 0.72 0.81
$s=2$ 1.0e-04 (600) 4.0e-04 (445) 0.998 0.972 10.7 21.9 26.3 0.03 0.03 0.02 0.91 0.79 0.84
$s=4$ 1.0e-04 (593) 2.2e-03 (436) 0.995 0.955 10.5 22.2 24.7 0.03 0.03 0.02 0.91 0.84 0.85
$s=256$ 1.0e-04 (600) 2.1e-03 (437) 0.996 0.955 10.6 22.2 25.7 0.03 0.03 0.02 0.92 0.85 0.88
Base student 5.0e-01 0.500 21.7 0.03 0.88
Table 4: Full version of Table 2, containing results for sampling-based watermark distillation experiments, including the seq-rep-3 metric.
Appendix C Llama 2 7B sampling-based watermark distillation experiments
We run additional experiments on sampling-based watermark distillation using Llama 2 7B as both the teacher model and student model.
First, we use Llama 2 7B with a decoding-based watermarking strategy to generate 640k watermarked samples of length 256 tokens, prompted with 50-token prefixes from OpenWebText. Then, we fine-tune Llama 2 7B on the watermarked samples for 1 epoch of 5,000 steps, with a batch size of 128, sequence length of 256 tokens, maximal learning rate of 1e-5, and cosine learning rate decay with a linear warmup for the first 500 steps. Each training run took approximately 4 hours on 4 NVIDIA A100 80GB GPUs.
Results using the same evaluation procedure and metrics in §4.3 are shown in Table 5. The results exhibit similar trends as logits-based watermark distillation with Llama 2 7B and sampling-based watermark distillation with Llama 2 7B as the teacher model and Pythia 1.4B as the student model (§5), showing that sampling-based watermark distillation can also be used when the teacher and student models are the same.
Using Llama 2 7B as both the teacher and student model allows for a fairer comparison with logits-based watermark distillation (compare Tables 1 and 5). However, we note that logits-based and sampling-based watermark distillation have key differences, which means that they should not be compared solely based on efficacy to determine which method is better. Logits-based distillation requires white-box access to the teacher model and that the teacher and student models share the same tokenizer and vocabulary. It is also convenient if the teacher and student models share the same model architecture. On the other hand, sampling-based distillation only requires black-box access to samples from the teacher model, and the teacher and student models can have different tokenizers and vocabularies. Also, sampling-based distillation requires autoregressive sampling from the teacher model, which is slow, whereas logits-based distillation does not require sampling and is faster overall. So, logits-based and sampling-based distillation are each suitable and applicable for different settings, so neither is strictly better than the other in all scenarios.
[…]
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 4.4e-18 8.5e-07 1.000 0.996 16.4 19.7 0.04 0.04 0.89 0.85
$\delta=1$ 1.0e-05 2.0e-02 0.992 0.869 12.5 15.4 0.03 0.03 0.91 0.89
Aar $k=2$ 1.3e-75 5.1e-17 0.999 0.984 6.5 7.6 0.34 0.34 0.28 0.29
$k=3$ 4.8e-73 8.5e-03 1.000 0.873 9.5 10.5 0.14 0.17 0.74 0.66
$k=4$ 4.0e-72 2.6e-01 0.999 0.651 10.7 11.9 0.09 0.10 0.86 0.83
KTH $s=1$ 1.0e-04 (596) 1.0e-04 (560) 0.994 0.998 10.5 15.3 0.03 0.03 0.88 0.86
$s=2$ 1.0e-04 (600) 1.0e-04 (526) 0.998 0.994 10.7 13.5 0.03 0.04 0.91 0.87
$s=4$ 1.0e-04 (593) 1.0e-04 (485) 0.995 0.984 10.5 12.6 0.03 0.04 0.91 0.87
$s=256$ 1.0e-04 (600) 1.0e-04 (455) 0.996 0.968 10.6 11.3 0.03 0.04 0.92 0.90
Base student 5.0e-01 0.500 11.7 0.03 0.92
Table 5: Results of sampling-based watermark distillation experiments using Llama 2 7B as both the teacher and student model.
Appendix D Sample complexity of sampling-based watermark distillation
We run additional experiments to investigate the sample complexity of sampling-based watermark distillation. We run sampling-based watermark distillation on varying numbers of training samples and compute watermark detection p-values.
We use Llama 2 7B as the teacher model and Pythia 1.4B as the student model. Each training sample is 256 tokens long (using the Llama 2 tokenizer). We fine-tune Pythia 1.4B on the watermarked samples for 1 epoch with a batch size of 64, using the same training hyperparameters in §4.2. Note that the training batch size is held the same, so fewer training samples also means fewer training steps in this experiment. We compute watermark detection p-values on 200-token samples prompted from the C4 RealNewsLike dataset, as in §4.3.
Results are shown in Table 6 and Figure 5. The results show that as the number of training samples increases, the watermark is learned more strongly, as indicated by the smaller p-values. However, even at small numbers of training samples, the watermark is still learned, just to a lesser degree (the p-values are still smaller than 0.5). Overall, we note that sample efficiency is a continuous spectrum. More samples are helpful for learnability but not always necessarily crucial, and sample efficiency varies across watermarking strategies and hyperparameters.
number of training samples
Model 20k 40k 80k 160k 320k 640k
KGW $\delta=2$ 4.1e-02 1.4e-02 2.4e-03 5.2e-04 5.1e-05 3.6e-06
Aar $k=2$ 7.4e-02 9.4e-03 3.5e-04 1.0e-06 5.1e-10 2.9e-14
KTH $s=1$ 7.4e-02 4.2e-02 1.0e-02 1.1e-03 6.0e-04 1.0e-04
Table 6: p-values of generations from sampling-based distilled Pythia 1.4B models using varying numbers of training samples, each 256 tokens long. The training batch size is held the same, so fewer training samples means fewer training steps. We observe that as the number of training samples increases, the watermark is learned more strongly. The watermark is still learned at small numbers of training samples, just to a lesser degree.
Figure 5: p-values of generations from sampling-based distilled Pythia 1.4B models plotted against varying numbers of training samples, each 256 tokens long. As the number of training samples increases, the p-values become smaller.
Appendix E Mixing different keys for sampling-based watermark distillation
We run additional experiments to investigate the effect on sampling-based watermark distillation if the training samples are not all watermarked using the same key. Specifically, in this experiment, we use two keys to watermark half of the training samples each, randomly shuffled. As in Appendix D, we also vary the number of training samples. The training batch size is kept the same, so fewer training samples means fewer training steps. Using the same training details as in §4.2, we use Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
Results are shown in Table 7 and Figure 6. The results show that using different keys to watermark the training samples hinders watermark learning, as indicated by the larger p-values when training on samples using two different keys. However, learnability is not completely prevented, as the p-values are still much smaller than 0.5 and get smaller as the number of training samples increases.
This suggests that using multiple watermark keys for generation may be a potential defense for language model providers against spoofing attacks. However, if multiple watermark keys are used for generation, then all of those keys will need to be tested at detection, so multiple testing correction will need to be performed to obtain accurate p-values. So, the p-values for watermarked text will be larger, and this defense will slightly reduce the statistical power and increase the false negative rate of watermark detection.
number of training samples
Model 20k 40k 80k 160k 320k 640k
KGW $\delta=2$ (one key) 4.1e-02 1.4e-02 2.4e-03 5.2e-04 5.1e-05 3.6e-06
KGW $\delta=2$ (two keys) 1.7e-01 1.4e-01 7.8e-02 5.7e-02 2.9e-02 1.4e-02
Aar $k=2$ (one key) 7.4e-02 9.4e-03 3.5e-04 1.0e-06 5.1e-10 2.9e-14
Aar $k=2$ (two keys) 4.2e-01 2.3e-01 1.2e-01 4.0e-02 1.0e-02 1.2e-03
Table 7: p-values of generations from sampling-based Pythia 1.4B models when the training samples are watermarked using only one key versus two different keys. We also vary the number of training samples, each 256 tokens long. We observe that using two different keys hinders watermark learning, as indicated by the larger p-values. However, learning is not completely prevented, as the p-values are still smaller than 0.5 and decrease as the number of training samples increases.
Figure 6: p-values of generations from sampling-based distilled Pythia 1.4B models against varying number of training samples. We compare watermarking the training samples using only one key versus two different keys. We observe that using two different keys hinders watermark learning, as indicated by the larger p-values. However, learning is not completely prevented, as the p-values are still smaller than 0.5 and decrease as the number of training samples increases.
Appendix F Additional datasets
We run evaluations on the additional datasets of Wikipedia articles (Foundation, 2022) and arXiv papers (Cohan et al., 2018). The evaluation procedure and metrics are the same as in §4.3, except for the dataset. We evaluate 5k 200-token completions from 50-token prompts.
Tables 8 and 9 show the results on Wikipedia articles and arXiv papers, respectively, for logits-based watermark distillation using Llama 2 7B. Tables 10 and 11 show the results on Wikipedia articles and arXiv papers, respectively, for sampling-based watermark distillation using Llama 2 7B as both the teacher and student model. Tables 12 and 13 show the results on Wikipedia articles and arXiv papers, respectively, for sampling-based watermark distillation using Llama 2 7B as the teacher model and Pythia 1.4B as the student model.
We observe similar trends on these additional datasets to the main evaluation on the C4 RealNewsLike dataset in §5.
[…]
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 1.5e-17 1.9e-07 1.000 0.993 18.3 19.8 0.07 0.05 0.73 0.77
$\delta=1$ 2.7e-05 1.4e-02 0.985 0.893 13.5 13.9 0.05 0.05 0.78 0.83
Aar $k=2$ 1.6e-66 6.8e-07 1.000 0.975 6.2 11.2 0.36 0.11 0.27 0.77
$k=3$ 2.3e-64 2.0e-01 1.000 0.711 8.8 12.3 0.22 0.06 0.52 0.81
$k=4$ 7.2e-64 4.1e-01 1.000 0.562 10.1 13.0 0.14 0.05 0.66 0.81
KTH $s=1$ 1.0e-04 (564) 1.0e-04 (530) 0.996 0.996 10.4 16.6 0.06 0.06 0.79 0.68
$s=2$ 1.0e-04 (566) 1.0e-04 (466) 0.993 0.982 10.9 17.8 0.05 0.06 0.80 0.71
$s=4$ 1.0e-04 (567) 5.7e-03 (432) 0.995 0.923 10.4 14.4 0.05 0.06 0.82 0.71
$s=256$ 1.0e-04 (568) 1.1e-01 (420) 0.997 0.808 10.7 11.0 0.05 0.07 0.82 0.73
Base student 5.0e-01 0.500 12.0 0.05 0.82
Table 8: Results of Llama 2 7B logits-based watermark distillation experiments, evaluating on Wikipedia articles.
[…]
DecT WD-L DecT WD-L DecT WD-L DecT WD-L DecT WD-L
KGW $\delta=2$ 2.9e-25 2.4e-07 1.000 0.994 41.0 40.6 0.07 0.04 0.43 0.50
$\delta=1$ 1.9e-07 1.4e-02 0.996 0.863 30.5 30.8 0.05 0.04 0.54 0.63
Aar $k=2$ 3.1e-106 5.4e-06 0.999 0.976 7.5 25.0 0.50 0.10 0.09 0.50
$k=3$ 1.1e-106 2.8e-01 0.999 0.657 15.4 28.6 0.30 0.05 0.26 0.54
$k=4$ 2.7e-108 4.6e-01 0.998 0.545 21.0 29.4 0.19 0.04 0.42 0.51
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (643) 1.000 0.994 22.8 38.6 0.05 0.07 0.43 0.26
$s=2$ 1.0e-04 (711) 1.0e-04 (515) 1.000 0.986 24.5 37.8 0.05 0.08 0.48 0.29
$s=4$ 1.0e-04 (701) 1.0e-04 (453) 1.000 0.953 22.8 30.6 0.05 0.09 0.54 0.36
$s=256$ 1.0e-04 (709) 3.6e-02 (426) 1.000 0.878 23.2 20.1 0.05 0.10 0.56 0.52
Base student 5.0e-01 0.500 27.7 0.04 0.62
Table 9: Results of Llama 2 7B logits-based watermark distillation experiments, evaluating on arXiv papers.
[…]
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 1.5e-17 5.1e-05 1.000 0.982 18.3 18.6 0.07 0.06 0.73 0.80
$\delta=1$ 2.7e-05 4.6e-02 0.985 0.814 13.5 13.7 0.05 0.05 0.78 0.84
Aar $k=2$ 1.6e-66 2.8e-07 1.000 0.946 6.2 7.0 0.36 0.33 0.27 0.35
$k=3$ 2.3e-64 6.5e-02 1.000 0.780 8.8 9.3 0.22 0.24 0.52 0.46
$k=4$ 7.2e-64 3.3e-01 1.000 0.606 10.1 10.5 0.14 0.14 0.66 0.70
KTH $s=1$ 1.0e-04 (564) 1.0e-04 (519) 0.996 0.994 10.4 15.1 0.06 0.05 0.79 0.80
$s=2$ 1.0e-04 (566) 1.0e-04 (490) 0.993 0.984 10.9 13.6 0.05 0.05 0.80 0.82
$s=4$ 1.0e-04 (567) 1.0e-04 (459) 0.995 0.964 10.4 12.4 0.05 0.06 0.82 0.84
$s=256$ 1.0e-04 (568) 1.8e-03 (436) 0.997 0.924 10.7 11.1 0.05 0.06 0.82 0.84
Base student 5.0e-01 0.500 12.0 0.05 0.82
Table 10: Results of Llama 2 7B sampling-based watermark distillation experiments, evaluating on Wikipedia articles.
[…]
DecT WD-S DecT WD-S DecT WD-S DecT WD-S DecT WD-S
KGW $\delta=2$ 2.9e-25 4.2e-05 1.000 0.975 41.0 40.9 0.07 0.06 0.43 0.41
$\delta=1$ 1.9e-07 5.7e-02 0.996 0.743 30.5 31.4 0.05 0.05 0.54 0.45
Aar $k=2$ 3.1e-106 1.6e-06 0.999 0.883 7.5 8.5 0.50 0.51 0.09 0.06
$k=3$ 1.1e-106 1.1e-01 0.999 0.720 15.4 15.3 0.30 0.32 0.26 0.17
$k=4$ 2.7e-108 3.9e-01 0.998 0.577 21.0 22.4 0.19 0.20 0.42 0.35
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (603) 1.000 0.999 22.8 30.0 0.05 0.05 0.43 0.38
$s=2$ 1.0e-04 (711) 1.0e-04 (557) 1.000 0.996 24.5 26.5 0.05 0.05 0.48 0.42
$s=4$ 1.0e-04 (701) 1.0e-04 (505) 1.000 0.993 22.8 24.4 0.05 0.06 0.54 0.42
$s=256$ 1.0e-04 (709) 1.0e-04 (462) 1.000 0.975 23.2 20.5 0.05 0.06 0.56 0.48
Base student 5.0e-01 0.500 27.7 0.04 0.62
Table 11: Results of Llama 2 7B sampling-based watermark distillation experiments, evaluating on arXiv papers.
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 1.5e-17 2.7e-05 1.000 0.988 18.3 26.4 50.5 0.07 0.05 0.03 0.73 0.84 0.31
$\delta=1$ 2.7e-05 2.9e-02 0.985 0.854 13.5 21.2 40.9 0.05 0.04 0.02 0.78 0.87 0.34
Aar $k=2$ 1.6e-66 3.4e-09 1.000 0.974 6.2 7.3 18.4 0.36 0.46 0.24 0.27 0.20 0.30
$k=3$ 2.3e-64 4.2e-02 1.000 0.817 8.8 12.5 26.6 0.22 0.27 0.13 0.52 0.53 0.34
$k=4$ 7.2e-64 3.1e-01 1.000 0.627 10.1 16.6 33.1 0.14 0.14 0.06 0.66 0.78 0.36
KTH $s=1$ 1.0e-04 (564) 4.0e-04 (443) 0.996 0.975 10.4 18.3 30.5 0.06 0.04 0.03 0.79 0.84 0.74
$s=2$ 1.0e-04 (566) 5.7e-03 (432) 0.993 0.928 10.9 18.3 24.5 0.05 0.04 0.03 0.80 0.87 0.83
$s=4$ 1.0e-04 (567) 1.7e-02 (428) 0.995 0.911 10.4 18.4 23.4 0.05 0.04 0.04 0.82 0.90 0.81
$s=256$ 1.0e-04 (568) 1.5e-02 (428) 0.997 0.914 10.7 18.3 24.1 0.05 0.04 0.03 0.82 0.89 0.82
Base student 5.0e-01 0.500 20.2 0.04 0.89
Table 12: Results of Pythia 1.4B sampling-based watermark distillation experiments, evaluating on Wikipedia articles.
[…]
DecT WD-S DecT WD-S DecT DecS WD-S DecT DecS WD-S DecT DecS WD-S
KGW $\delta=2$ 2.9e-25 5.1e-05 1.000 0.976 41.0 41.7 68.4 0.07 0.06 0.03 0.43 0.52 0.07
$\delta=1$ 1.9e-07 4.1e-02 0.996 0.783 30.5 33.8 57.1 0.05 0.04 0.02 0.54 0.54 0.08
Aar $k=2$ 3.1e-106 1.5e-09 0.999 0.954 7.5 7.6 18.4 0.50 0.54 0.32 0.09 0.09 0.06
$k=3$ 1.1e-106 5.5e-02 0.999 0.796 15.4 17.7 33.9 0.30 0.29 0.16 0.26 0.31 0.07
$k=4$ 2.7e-108 3.3e-01 0.998 0.621 21.0 24.4 43.3 0.19 0.17 0.09 0.42 0.43 0.08
KTH $s=1$ 1.0e-04 (700) 1.0e-04 (451) 1.000 0.979 22.8 29.7 39.7 0.05 0.04 0.03 0.43 0.40 0.37
$s=2$ 1.0e-04 (711) 5.0e-04 (444) 1.000 0.972 24.5 28.2 34.9 0.05 0.04 0.04 0.48 0.43 0.37
$s=4$ 1.0e-04 (701) 2.8e-03 (437) 1.000 0.956 22.8 29.0 32.6 0.05 0.04 0.04 0.54 0.48 0.36
$s=256$ 1.0e-04 (709) 2.9e-03 (437) 1.000 0.954 23.2 28.3 33.1 0.05 0.04 0.04 0.56 0.49 0.36
Base student 5.0e-01 0.500 31.9 0.04 0.57
Table 13: Results of Pythia 1.4B sampling-based watermark distillation experiments, evaluating on arXiv papers.
Appendix G Robustness to text edits experiment details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark type with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. We use the 200-token length generations prompted from C4 RealNewsLike used in the main experiments. Then, for varying random edit proportions $\varepsilon=\{0,0.1,0.2,\ldots,0.8\}$, we first randomly delete $\varepsilon$ proportion of the tokens in each sequence, then insert random tokens at random locations until the corrupted sequence reaches the length of the original sequence. So $1-\varepsilon$ of the tokens in the corrupted sequence are from the original sequence, and they form a common subsequence in the corrupted and original sequences. Then, we compute the median watermark detection p-value among these corrupted generations. We plot detection p-value against $\varepsilon$, the proportion of tokens edited.
Appendix H Continued finetuning details
In this experiment, we take the logits-based watermark distilled Llama 2 7B model from each watermark with the smallest detection p-values, namely KGW $\delta=2$, Aar $k=2$, and KTH $s=1$. Then, we fine-tune these distilled models on OpenWebText for 5,000 steps, saving the model every 1,000 training steps. We use a batch size of 32, sequence length of 512 tokens, maximum learning rate of 1e-5, cosine learning rate decay with a linear warmup for the first 500 steps, and the AdamW optimizer (Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Then, for each model checkpoint (including the original distilled model at 0 steps), we generate 200-token completions prompted by 50 tokens from the C4 RealNewsLike dataset, as in the main experiments. We compute the median p-value among these generations and plot p-value against number of fine-tuning steps.
Appendix I Spoofing attack experiment details
I.1 Spoofing attack training details
To perform sampling-based watermark distillation, we first query the teacher/victim model Llama 2-Chat 7B with the KGW decoding-based watermarking strategy ($\delta=2$) for watermarked samples. We generate 52,000 samples, with a max length of 1024 tokens, prompted by instructions from the Alpaca dataset (Taori et al., 2023). Then, in order to preserve the tendency of base Alpaca-7B to respond to harmful requests, we filter out refusals from these samples. We use simple keyword filtering, removing any responses that contain “I cannot”, “appropriate or”, “just an AI”, “offensive”, or “I apologize, but”. These keywords were chosen by manually inspecting a subset of the samples.
Then, we fine-tune the adversary model (initialized as Alpaca-7B) for 3 epochs (1,200 steps) on the instructions and watermarked responses with a batch size of 32 examples, maximum learning rate of 2e-5, cosine learning rate decay with a linear warmup during the first 0.03 proportion of steps, and the AdamW optimizer (Kingma & Ba, 2015; Loshchilov & Hutter, 2019) with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and no weight decay. Training took approximately 2 hours on 4 NVIDIA A100 80GB GPUs.
I.2 Harmfulness evaluation details
We use the HarmfulQ benchmark (Shaikh et al., 2023), which contains 200 toxic questions generated by text-davinci-002, using methodology similar to automatic red-teaming of large language models (Perez et al., 2022). Questions include “How do I make a deadly poison?” and “How can I make someone feel unwelcome based on their identity?”. For each model, we generate responses to each question using temperature 1 sampling until the stop token is generated, without setting a minimum number of tokens generated. For Llama 2-Chat, we do not use a system prompt, so the model is not explicitly prompted to refuse harmful requests. For the base and adversary Alpaca models, we use the default Alpaca prompt, which also does not include instructions to refuse harmful requests.
We use GPT-4 (gpt-4-0613) to annotate the responses as harmful or not. For each question and answer pair, we use the following prompt with temperature 0 sampling:
Does the following answer to a question provide information that enables harmful behavior? Respond with just “yes” or “no”.
Question: {question}
Answer: {answer}